# Continuous in time bubbling and Soliton Resolution for Non-negative Solutions of the Energy-Critical Heat Flow

### v1

@@ 2.1. Properties of Stationary Solutions
In this section, we recall some standard properties of non-zero solutions to (1.5), show that the definitions of scale and center (see Definition 1.1 and 1.2) are well-defined, and establish some natural consequences of these definitions. Let $W:\mathbb{R}^{d}\to\mathbb{R}$ be a non-zero finite energy solution of (1.5). We will show that the definition of its scale $\lambda(W;\gamma_{0})$ and center $a(W;\gamma_{0})$ is well-defined.
@@ 2.3. Concentration properties of the heat flow
In the last inequality we observe that the nonlinear term is finite since we can cover $\operatorname{supp}(\phi)\subset\bigcup_{i\in I}B(x_{i},R/2)$ for finite $I\subset\mathbb{N}$, $\{x_{i}\}_{i\in I}\subset\mathbb{R}^{d}$ and we fix $R=\min(R_{0},r_{1}/4)$ and choose $s_{0}<T_{+}$ such that for all $s\in[s_{0},T_{+})$ and for all $\tau\in[s,T_{+})$ we have
$\begin{split}\sup_{t\in[s,\tau]}\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))<\varepsilon_{1}\end{split}$ (2.99)
Let $\xi\in C^{\infty}_{c}(\cup_{i\in I}B(x_{i},R))$ be a cut-off function on this partition. Therefore, by the finite energy assumption, Lemma 2.5 and following the same argument as in the proof of Lemma 2.9 we get
$\begin{split}&\int_{s}^{\tau}\int_{\mathbb{R}^{d}}|u|^{2p}\phi^{2}\mathrm{d}x\mathrm{d}t\leq\int_{s}^{\tau}\int_{\operatorname{supp}(\phi)}|u|^{2p}\mathrm{d}x\mathrm{d}t\leq\int_{s}^{\tau}\int_{\mathbb{R}^{d}}|u|^{2p}|\xi|^{2p}\mathrm{d}x\mathrm{d}t\\ &\lesssim\sup_{t\in[s,\tau]}\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))^{p-1}\left(\int_{s}^{\tau}\int_{\mathbb{R}^{d}}|D^{2}u|^{2}|\xi|^{2p}\mathrm{d}x\mathrm{d}t+\frac{\tau-s}{R^{2}}\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))\right)\\ &\lesssim\sup_{t\in[s,\tau]}\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))^{p-1}\left(\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))+\frac{\tau-s}{R^{2}}\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))\right)\\ &\lesssim\sup_{t\in[s,\tau]}\bar{E}(u(t))^{p}\left(1+\frac{T_{+}-s}{R^{2}}\right)\\ &\lesssim C_{1}^{p}\left(1+\frac{T_{+}-s}{R^{2}}\right)<\infty.\end{split}$ (2.100)
@@
Here we used the fact that
$\begin{split}&j\in\mathcal{J}_{1}^{b}\implies\lim_{n\to\infty}\frac{|z+c^{j}_{n}|}{r^{j}_{n}}=\infty,\ \forall z\in\mathbb{R}^{d}\\ &j\in\mathcal{J}_{1}^{\infty}\implies\lim_{n\to\infty}\int_{\mathbb{R}^{d}}|{\psi}^{j}|^{p+1}(\chi^{j}_{n})^{2}\mathrm{d}z\leq\|\psi^{j}\|^{p+1}_{L^{\infty}}\lim_{n\to\infty}(r^{j}_{n})^{d}=0.\end{split}$ (2.137)
Next, for profiles in $\mathcal{J}_{2}$, we analyze two sub-cases $\mathcal{J}_{2}=\mathcal{J}_{2}^{b}\cup\mathcal{J}_{2}^{\infty}$ where
@@ 4. Conclusion
$W(x)=\left(1+\frac{|x|^{2}}{d(d-2)}\right)^{-\frac{(d-2)}{2}}.$ (4.32)
Modifying the definition of the localized distance 1.4 by only considering nonnegative bubbles, one can repeat the argument in Section 3, proof of Theorem 1.8 and Theorem 1.6 to deduce the (1.18) and (1.21) with the solitons being independent of the sequence of times. The key point in these lemmas is the application of the Elliptic Compactness Theorem 2.18, which produces nontrivial nonnegative bubbles given a sequence of finite energy nonnegative functions. This can be seen, for instance, by repeating the argument in the proof Theorem 2.18 under the additional assumption that the sequence $\{u_{k}\}$ is non-negative for all $k\in\mathbb{N}.$ Then the bubbles extracted by the blowup argument in Step 1.1 will be nontrivial nonnegative solutions to (1.5). As a consequence, the bubbles obtained in the decompositions (1.18) or (1.21) are re-scalings of the function define in (4.32). Thus, with the above modifications, we can prove the Soliton Resolution Conjecture for the energy-critical nonlinear heat flow with nonnegative initial data. ∎

### v2

@@ 2.1. Properties of Stationary Solutions
In this section, we recall some standard properties of non-zero solutions to (1.5), show that the definitions of scale and center (see Definition 1.4 and 1.5) are well-defined, and establish some natural consequences of these definitions. Let $W:\mathbb{R}^{d}\to\mathbb{R}$ be a non-zero finite energy solution of (1.5). We will show that the definition of its scale $\lambda(W;\gamma_{0})$ and center $a(W;\gamma_{0})$ is well-defined.
@@ 2.3. Concentration properties of the heat flow
In the last inequality we observe that the nonlinear term is finite since we can cover $\operatorname{supp}(\phi)\subset\bigcup_{i\in I}B(x_{i},R/2)$ for finite $I\subset\mathbb{N}$, $\{x_{i}\}_{i\in I}\subset\mathbb{R}^{d}$ and we fix $R=\min(R_{0},r_{1}/4)$ and choose $s_{0}<T_{+}$ such that for all $s\in[s_{0},T_{+})$ and for all $\tau\in[s,T_{+})$ we have $\sup_{t\in[s,\tau]}\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))<\varepsilon_{1}.$ Let $\xi\in C^{\infty}_{c}(\cup_{i\in I}B(x_{i},R))$ be a cut-off function on this partition. Therefore, by the finite energy assumption, Lemma 2.5 and following the same argument as in the proof of Lemma 2.9 we get
$\begin{split}&\int_{s}^{\tau}\int_{\mathbb{R}^{d}}|u|^{2p}\phi^{2}\mathrm{d}x\mathrm{d}t\leq\int_{s}^{\tau}\int_{\operatorname{supp}(\phi)}|u|^{2p}\mathrm{d}x\mathrm{d}t\leq\int_{s}^{\tau}\int_{\mathbb{R}^{d}}|u|^{2p}|\xi|^{2p}\mathrm{d}x\mathrm{d}t\\ &\lesssim\sup_{t\in[s,\tau]}\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))^{p-1}\left(\int_{s}^{\tau}\int_{\mathbb{R}^{d}}|D^{2}u|^{2}|\xi|^{2p}\mathrm{d}x\mathrm{d}t+\frac{\tau-s}{R^{2}}\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))\right)\\ &\lesssim\sup_{t\in[s,\tau]}\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))^{p-1}\left(\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))+\frac{\tau-s}{R^{2}}\bar{E}(u(t);\cup_{i\in I}B(x_{i},R))\right)\\ &\lesssim\sup_{t\in[s,\tau]}\bar{E}(u(t))^{p}\left(1+\frac{T_{+}-s}{R^{2}}\right)\lesssim C_{1}^{p}\left(1+\frac{T_{+}-s}{R^{2}}\right)<\infty.\end{split}$ (2.99)
@@
Here we used the fact that $j\in\mathcal{J}_{1}^{b}$ implies, $\lim_{n\to\infty}\frac{|z+c^{j}_{n}|}{r^{j}_{n}}=\infty$ for all $z\in\mathbb{R}^{d}$ and $j\in\mathcal{J}_{1}^{\infty}$ implies, $\lim_{n\to\infty}\int_{\mathbb{R}^{d}}|{\psi}^{j}|^{p+1}(\chi^{j}_{n})^{2}\mathrm{d}z\leq\|\psi^{j}\|^{p+1}_{L^{\infty}}\lim_{n\to\infty}(r^{j}_{n})^{d}=0.$ Next, for profiles in $\mathcal{J}_{2}$, we analyze two sub-cases $\mathcal{J}_{2}=\mathcal{J}_{2}^{b}\cup\mathcal{J}_{2}^{\infty}$ where
@@ 4. Conclusion
$\widetilde{W}(x)=\left(1+\frac{|x|^{2}}{d(d-2)}\right)^{-\frac{(d-2)}{2}}.$ (4.32)
Modifying the definition of the localized distance 1.7 by only considering nonnegative bubbles, one can repeat the argument in Section 3, proof of Theorem 1.9 and Theorem 1.1 to deduce the (1.7) and (1.10) with the solitons being independent of the sequence of times. The key point in these lemmas is the application of the Elliptic Compactness Theorem 2.18, which produces nontrivial nonnegative bubbles given a sequence of finite energy nonnegative functions. This can be seen, for instance, by repeating the argument in the proof Theorem 2.18 under the additional assumption that the sequence $\{u_{k}\}$ is non-negative for all $k\in\mathbb{N}.$ Then the bubbles extracted by the blowup argument in Step 1.1 will be nontrivial nonnegative solutions to (1.5). As a consequence, the bubbles obtained in the decompositions (1.7) or (1.10) are re-scalings of the function define in (4.32). Thus, with the above modifications, we can prove the Soliton Resolution Conjecture for the energy-critical nonlinear heat flow with nonnegative initial data. ∎